# Dual Quaternion SE(3) Synchronization with Recovery Guarantees

### v1

@@ Abstract
footnotetext: Equal contribution.
@@ 1 Introduction
$\mathrm{SO}(2)$ $z_{i}\leftarrow z_{i}/|z_{i}|$ $\mathrm{Iso}(E)^{*}\cong\mathrm{SO}(2)\Rightarrow$ No gap
@@
$\mathrm{SE}(3)$ Multi-step heuristic $\mathrm{Iso}(E)$ compact vs. non-compact $\mathrm{SE}(3)\Rightarrow$ Large mismatch
@@
However, matrix-based representations suffer from a fundamental limitation on the representation gap. Ideally, a group representation should align the isometry group of the nonzero eigenspace with the intrinsic global gauge symmetry of the synchronization problem. This alignment is essentially perfect for $\mathrm{SO}(2)$ under the unit complex representation: the eigenspace isometry matches $\mathrm{SO}(2)$, making rounding a trivial normalization. For $\mathrm{SO}(d)$, the eigenspace isometry expands to $\mathrm{O}(d)$, introducing only a mild gap (a reflection ambiguity) that is rigorously resolved via SVD projection. In contrast, as summarized in Table 1, the conventional $4\times 4$ matrix representation of $\mathrm{SE}(3)$ breaks this favorable correspondence. While the relaxed eigenspace possesses a compact orthogonal geometry, $\mathrm{SE}(3)=\mathrm{SO}(3)\ltimes\mathbb{R}^{3}$ is inherently non-compact. Consequently, the relaxed solution often lies far from the manifold, creating a large eigenspace-sync gap. Rounding is no longer a benign projection that merely corrects noise; instead, it must artificially enforce a group structure absent from the eigenspace geometry. This necessitates complex, multi-step heuristic procedures that are often unstable (Arrigoni et al., 2016b; Rosen et al., 2019). Although alternative representations exist, such as axis-angle pairs (Thunberg et al., 2014; Jin et al., 2020), augmented unit quaternions (Grisetti et al., 2010; Chen et al., 2024), and dual quaternions (Cheng et al., 2016; Srivatsan et al., 2016; Hadi et al., 2024) exist, they typically lack a unified algebraic framework that is both compact and amenable to rigorous optimization guarantees.
@@
$\mathop{\arg\min}\limits_{x\in\mathrm{UDQ}^{n}}\sum_{1\leq i<j\leq n}\left|C_{ij}-x_{i}x_{j}^{*}\right|^{2},$ (1.1)
@@
To efficiently tackle the nonconvex optimization problem (1.1), we build on the generalized power method (GPM) and develop a two-stage framework that combines a provably accurate spectral initializer with an iterative refinement scheme that preserves feasibility at every step. Variants of this spectral initializer with an iterative refinement strategy have proved effective for phase, orthogonal, and rotation group synchronization problems (Boumal, 2016; Ling, 2022; Liu et al., 2023; Zhu et al., 2023). However, establishing rigorous guarantees in the dual quaternion setting is substantially more challenging. First, dual quaternions form a ring with zero divisors rather than a field, so division is not generally well defined. Second, as the natural modulus on dual quaternions is merely a dual-number valued seminorm, a bounded modulus does not imply bounded components. This allows translations to grow arbitrarily large despite a finite modulus, thereby posing challenges for stability and convergence analysis. Third, identifying a dominant eigenpair typically invokes the lexicographic order on dual numbers, whereas estimation and convergence are naturally quantified in Euclidean metrics. By addressing these theoretical hurdles, we provide the first recovery guarantees for synchronization over the dual quaternion. Our main contributions are as follows:
@@ 1.1 Notation
Throughout the paper, the sets of dual numbers, quaternions, and dual quaternions are denoted by $\mathbb{D}$, $\mathbb{H}$, and $\mathbb{DH}$, respectively. For a dual number $a=a_{st}+a_{\mathcal{I}}\epsilon$ with $a_{st}\neq 0$, its inverse is $a^{-1}=a_{st}^{-1}\bigl(1-a_{\mathcal{I}}a_{st}^{-1}\epsilon\bigr)$. The operator $(\cdot)^{*}$ denotes the (dual) quaternion conjugate when applied to a scalar element, and the conjugate transpose when applied to a vector or matrix. The scalar part of a quaternion $q$ is denoted by $\mathrm{sc}(q)=\tfrac{1}{2}(q+q^{*})$. For two $n$-dimensional (dual) quaternion vectors $\bm{x}=(x_{1},\ldots,x_{n})$ and $\bm{y}=(y_{1},\ldots,y_{n})$, their inner product is defined as $\bm{x}^{*}\bm{y}=\sum_{i=1}^{n}x_{i}^{*}y_{i}$. The Frobenius norm of a (dual) quaternion matrix is denoted by $\|\cdot\|_{F}$, which reduces to the $2$-norm $\|\cdot\|_{2}$ for a (dual) quaternion vector. The operator norm of a (dual) quaternion matrix is denoted by $\|\cdot\|_{\mathrm{op}}$. The set of unit dual quaternions is denoted by $\mathrm{UDQ}:=\{q\in\mathbb{DH}:qq^{*}=1\}$. We write $\mathrm{UDQ}^{n}:=\{\bm{x}=(x_{1},\dots,x_{n}):x_{i}\in\mathrm{UDQ}\}$, that is, the set of $n$-dimensional vectors whose entries are unit dual quaternions. For $\bm{x}\in\mathbb{DH}^{n}$, the element-wise projection onto $\mathrm{UDQ}^{n}$ is defined by $\Pi(\bm{x})=\mathop{\operatorname{argmin}}_{\bm{u}\in\mathrm{UDQ}^{n}}\|\bm{u}-\bm{x}\|_{2}^{2}.$ For any $\bm{x},\bm{y}\in\mathbb{DH}^{n}$, the distance between them, defined up to a global unit dual quaternion alignment, is given by $\operatorname{d}(\bm{x},\bm{y})=\min_{z\in\mathrm{UDQ}}\|\bm{x}-\bm{y}z\|_{2}.$ We denote by $\operatorname{d}_{st}(\bm{x},\bm{y})$ and $\operatorname{d}_{\mathcal{I}}(\bm{x},\bm{y})$ the standard part and the imaginary part of $\operatorname{d}(\bm{x},\bm{y})$, respectively. For $z\in\mathbb{DH}$, we denote the normalization of $z$ onto $\mathrm{UDQ}$ as $\mathcal{N}(z)$. Given $\{a_{k}\}\subset\mathbb{D}$ and $\{c_{k}\}\subset\mathbb{R}$, we define $a_{k}=O_{\mathbb{D}}(c_{k})$ iff $(a_{k})_{st}=O(c_{k})$ and $(a_{k})_{\mathcal{I}}=O(c_{k})$.
@@ 2 Spectral Estimator
The least-squares Problem (1.1) can be reformulated as an equivalent quadratic program with quadratic constraints (QPQC) over unit dual quaternions:
$\min_{\bm{x}\in\mathrm{UDQ}^{n}}\ \left\|\bm{C}-\bm{x}\bm{x}^{*}\right\|_{F}^{2}.$ (2.1)
@@
For a Hermitian dual quaternion matrix $\bm{C}\in\mathbb{DH}^{n\times n}$, Problem (2.1) is equivalent to
@@
$\mathop{\arg\max}\limits_{\bm{x}\in\mathbb{DH}^{n},\ \|\bm{x}\|^{2}_{2}=n}\ \bm{x}^{*}\bm{C}\bm{x}$ (2.2)
@@
By Fact 2.2, the optimum of (2.2) is attained at the dominant right eigenvector $\bm{u}_{1}$ of $\bm{C}$, normalized so that $\|\bm{u}_{1}\|_{2}^{2}=n$. Recently, Cui and Qi (2024) proposed a power iteration method for computing the dominant eigenvalue and a corresponding eigenvector of a dual quaternion Hermitian matrix, and established linear convergence of the resulting iterates to the dominant eigenpair. The procedure is summarized in Algorithm 1.
@@
1:  Input: Hermitian matrix $\bm{C}\in\mathbb{DH}^{n\times n}$ with dominant right eigenvalue $\lambda_{1}$ satisfying $\lambda_{1,\mathrm{st}}\neq 0$, random initial point $\bm{w}^{0}\in\mathbb{DH}^{n}$ such that $(\bm{w}^{0}_{st})^{*}\bm{u}_{1,st}\neq 0$.
@@
The mild assumption $\lambda_{1,st}\neq 0$ and $(\bm{w}^{0}_{st})^{*}\bm{u}_{1,st}\neq 0$ ensures that dominant eigenspace exists and the initialization has a nonzero standard part projection onto the dominant eigenspace. If $(\bm{w}^{0}_{st})^{*}\bm{u}_{1,st}=0$, then the power iteration cannot amplify (and hence cannot converge to) the dominant eigenspace. Under these assumptions, the standard part iterate remains nonzero, i.e., $\bm{y}^{k}_{\mathrm{st}}\neq\bm{0}$, making the inverse well-defined.
@@
We first define the normalization operator $\mathcal{N}(\cdot)$, which maps a dual quaternion $x\in\mathbb{DH}$ to a unit dual quaternion $\mathcal{N}(x)\in\mathrm{UDQ}$. Let $u:=\mathcal{N}(x)$. If $x_{st}\neq 0$, then
$u_{st}=\frac{x_{st}}{|x_{st}|},\ u_{\mathcal{I}}=\frac{x_{\mathcal{I}}}{|x_{st}|}-\frac{x_{st}}{|x_{st}|}\mathrm{sc}\left(\frac{x_{st}^{*}}{|x_{st}|}\frac{x_{\mathcal{I}}}{|x_{st}|}\right).$
If $x_{st}=0$ and $x_{\mathcal{I}}\neq 0$, then
$u_{st}=\frac{x_{\mathcal{I}}}{|x_{\mathcal{I}}|},\ u_{\mathcal{I}}\text{ is any quaternion satisfying }\mathrm{sc}(x_{\mathcal{I}}^{*}u_{\mathcal{I}})=0.$
@@
$\widetilde{\bm{y}}=\begin{cases}\mathcal{N}(y_{i})&\text{if }y_{i}\neq 0,\\ \mathcal{N}(\bm{e}^{*}\bm{y})&\text{otherwise},\end{cases}$
where $\bm{e}\in\mathbb{DH}^{n}$ is any vector satisfying $\bm{e}^{*}\bm{y}\neq 0$. If $\bm{y}=\bm{0}$, we set $\widetilde{\bm{y}}=\bm{1}^{n}$. Then $\tilde{\bm{y}}$ is the projection of $\bm{y}$ onto $\mathrm{UDQ}^{n}$, i.e., $\widetilde{\bm{y}}=\Pi(\bm{y})$.
@@
$\widetilde{\bm{x}}:=\Pi(\bm{u}_{1}),$ (2.3)
which lies in $\mathrm{UDQ}^{n}$ by construction. We next show that $\widetilde{\bm{x}}$ remains close to the ground-truth signal $\hat{\bm{x}}$.
@@
For $\widetilde{\bm{x}}$ given by (2.3), we have
@@
Theorem 2.8 follows from the Lipschitz property of the normalization/projection map in Lemma 2.5. In particular, it implies that $\widetilde{\bm{x}}$ achieves the same order of accuracy as the (generally infeasible) spectral solution $\bm{u}_{1}$ to (2.2), while ensuring feasibility in $\mathrm{UDQ}^{n}$.
@@ 3 Iterative Estimation Performance
Although Proposition 2.4 and Theorem 2.8 show that both the (infeasible) spectral solution of (2.2) and its rounded version yield accurate estimates close to the ground-truth signal, these estimators alone do not directly solve (2.1) beyond serving as a starting point. To obtain a fully iterative refinement procedure whose iterates remain feasible, we develop a dual quaternion generalized power method (DQGPM) as the second-stage algorithm.
@@
1:  Input: Hermitian matrix $\bm{C}\in\mathbb{DH}^{n\times n}$, initializer $\bm{x}^{0}=\bm{u}_{1}\in\mathbb{DH}^{n}$.
@@
We need to distinguish the algebraic order used in Section 2 from the error metric used in our convergence analysis. While the spectral estimator relies on the lexicographical order of dual numbers to define the dominant eigenpair (strictly prioritizing the standard part), SE(3) synchronization requires bounded errors in both rotation and translation. Therefore, in this section, we analyze the estimation error by controlling the real-valued magnitudes of the standard component $d_{st}$ and the dual component $d_{\mathcal{I}}$ individually. This component-wise analysis is consistent with the algebraic structure of dual numbers, where the convergence of the dual part ($d_{\mathcal{I}}$) typically depends on the accuracy of the standard part ($d_{st}$).
@@
Suppose that the measurement noise satisfies $\|\bm{\Delta}\|_{\mathrm{op},st}\leq n/350$, $\|\bm{\Delta}\|_{\mathrm{op},\mathcal{I}}\leq n/300$, and the initialization is given by $\bm{x}^{0}=\bm{u}_{1}$. Then the sequence generated by Algorithm 2 satisfies
[…]
and
$\operatorname{d}_{\mathcal{I}}(\bm{x}^{k},\hat{\bm{x}})\leq\left(\frac{1}{10}\right)^{k}\operatorname{d}_{\mathcal{I}}(\bm{x}^{0},\hat{\bm{x}})+k\left(\frac{1}{10}\right)^{k-1}\frac{1}{8}\operatorname{d}_{st}(\bm{x}^{0},\hat{\bm{x}})+\frac{700}{53n}\|\bm{\Delta}\hat{\bm{x}}\|_{2,\mathcal{I}}+\frac{875}{477n}\|\bm{\Delta}\hat{\bm{x}}\|_{2,st}.$ (3.2)

### v2

@@ Abstract
SE(3) Synchronization, Dual Quaternions, Spectral Methods
@@ 1 Introduction
$\mathrm{SO}(2)$ $z_{i}\leftarrow z_{i}/|z_{i}|$ $\mathrm{Iso}(E)\cong\mathrm{SO}(2)\Rightarrow$ No gap
@@
$\mathrm{SE}(3)$ Multi-step heuristic $\mathrm{Iso}(E)$ cannot encode translations $\Rightarrow$ Substantial mismatch
@@
However, matrix-based representations suffer from a fundamental limitation on the representation gap. Ideally, a group representation should align the isometry group of the nonzero eigenspace with the intrinsic global gauge symmetry of the synchronization problem. This alignment is essentially perfect for $\mathrm{SO}(2)$ under the unit complex representation: the eigenspace isometry matches $\mathrm{SO}(2)$, making rounding a trivial normalization. For $\mathrm{SO}(d)$, the eigenspace isometry expands to $\mathrm{O}(d)$, introducing only a mild gap (a reflection ambiguity) that is rigorously resolved via SVD projection. In contrast, as summarized in Table 1, the conventional $4\times 4$ matrix representation of $\mathrm{SE}(3)$ breaks this favorable correspondence. While the relaxed eigenspace possesses a compact orthogonal geometry, $\mathrm{SE}(3)=\mathrm{SO}(3)\ltimes\mathbb{R}^{3}$ is inherently non-compact. Consequently, the relaxed solution often lies far from the manifold, creating a large eigenspace-sync gap. Rounding is no longer a benign projection that merely corrects noise; instead, it must artificially enforce a group structure absent from the eigenspace geometry. This necessitates complex, multi-step heuristic procedures that are often unstable (Arrigoni et al., 2016b; Rosen et al., 2019). Although alternative representations exist, such as axis-angle pairs (Thunberg et al., 2014; Jin et al., 2020), augmented unit quaternions (Grisetti et al., 2010; Chen et al., 2025), and dual quaternions (Cheng et al., 2016; Srivatsan et al., 2016; Hadi et al., 2024), they typically lack a unified algebraic framework that is amenable to rigorous optimization guarantees. Concurrent with our work, Zhao et al. (2026) further studied matrix-based spectral estimators, introducing an additional anchoring step to resolve the right-orthogonal ambiguity of the eigenspace.
@@
$\mathop{\arg\min}\limits_{x\in\mathrm{UDQ}^{n}}\sum_{1\leq i<j\leq n}\left|C_{ij}-x_{i}x_{j}^{*}\right|^{2},$ (1)
@@
To efficiently tackle the nonconvex optimization problem (1), we build on the generalized power method (GPM) and develop a two-stage framework that combines a provably accurate spectral initializer with an iterative refinement scheme that preserves feasibility at every step. Variants of this spectral initializer with an iterative refinement strategy have proved effective for phase, orthogonal, and rotation group synchronization problems (Boumal, 2016; Ling, 2022; Liu et al., 2023; Zhu et al., 2023). However, establishing rigorous guarantees in the dual quaternion setting is substantially more challenging. First, dual quaternions form a ring with zero divisors rather than a field, so division is not generally well defined. Second, as the natural modulus on dual quaternions is merely a dual-number valued seminorm, a bounded modulus does not imply bounded components. This allows translations to grow arbitrarily large despite a finite modulus, thereby posing challenges for stability and convergence analysis. Third, identifying a dominant eigenpair typically invokes the lexicographic order on dual numbers, whereas estimation and convergence are naturally quantified in Euclidean metrics. By addressing these theoretical hurdles, we provide the first recovery guarantees for synchronization over the dual quaternion. Our main contributions are as follows:
@@ 1.1 Notation
Throughout the paper, the sets of dual numbers, quaternions, and dual quaternions are denoted by $\mathbb{D}$, $\mathbb{H}$, and $\mathbb{DH}$, respectively. For a dual number $a=a_{\mathrm{st}}+a_{\mathcal{I}}\epsilon$ with $a_{\mathrm{st}}\neq 0$, its inverse is $a^{-1}=a_{\mathrm{st}}^{-1}\bigl(1-a_{\mathcal{I}}a_{\mathrm{st}}^{-1}\epsilon\bigr)$. The operator $(\cdot)^{*}$ denotes the (dual) quaternion conjugate when applied to a scalar element, and the conjugate transpose when applied to a vector or matrix. The scalar part of a quaternion $q$ is denoted by $\mathrm{sc}(q)=\tfrac{1}{2}(q+q^{*})$. For two $n$-dimensional (dual) quaternion vectors $\bm{x}=(x_{1},\ldots,x_{n})$ and $\bm{y}=(y_{1},\ldots,y_{n})$, their inner product is defined as $\bm{x}^{*}\bm{y}=\sum_{i=1}^{n}x_{i}^{*}y_{i}$. The Frobenius norm of a (dual) quaternion matrix is denoted by $\|\cdot\|_{F}$, which reduces to the $2$-norm $\|\cdot\|_{2}$ for a (dual) quaternion vector. The operator norm of a (dual) quaternion matrix is denoted by $\|\cdot\|_{\mathrm{op}}$. The set of unit dual quaternions is denoted by $\mathrm{UDQ}:=\{q\in\mathbb{DH}:qq^{*}=1\}$. We write $\mathrm{UDQ}^{n}:=\{\bm{x}=(x_{1},\dots,x_{n}):x_{i}\in\mathrm{UDQ}\}$, that is, the set of $n$-dimensional vectors whose entries are unit dual quaternions. For $\bm{x}\in\mathbb{DH}^{n}$, the element-wise projection onto $\mathrm{UDQ}^{n}$ is defined by $\Pi(\bm{x})=\mathop{\operatorname{argmin}}_{\bm{u}\in\mathrm{UDQ}^{n}}\|\bm{u}-\bm{x}\|_{2}^{2}.$ For any $\bm{x},\bm{y}\in\mathbb{DH}^{n}$, the distance between them, defined up to a global unit dual quaternion alignment, is given by $\operatorname{d}(\bm{x},\bm{y})=\min_{z\in\mathrm{UDQ}}\|\bm{x}-\bm{y}z\|_{2}.$ We denote by $\operatorname{d}_{\mathrm{st}}(\bm{x},\bm{y})$ and $\operatorname{d}_{\mathcal{I}}(\bm{x},\bm{y})$ the standard part and the imaginary part of $\operatorname{d}(\bm{x},\bm{y})$, respectively. For $z\in\mathbb{DH}$, we denote the normalization of $z$ onto $\mathrm{UDQ}$ as $\mathcal{N}(z)$. Given $\{a_{k}\}\subset\mathbb{D}$ and $\{c_{k}\}\subset\mathbb{R}$, we define $a_{k}=O_{\mathbb{D}}(c_{k})$ iff $(a_{k})_{\mathrm{st}}=O(c_{k})$ and $(a_{k})_{\mathcal{I}}=O(c_{k})$.
@@ 2 Spectral Estimator
The least-squares Problem (1) can be reformulated as an equivalent quadratic program with quadratic constraints (QPQC) over unit dual quaternions:
$\min_{\bm{x}\in\mathrm{UDQ}^{n}}\ \left\|\bm{C}-\bm{x}\bm{x}^{*}\right\|_{F}^{2}.$ (3)
@@
For a Hermitian dual quaternion matrix $\bm{C}\in\mathbb{DH}^{n\times n}$, Problem (3) is equivalent to
@@
$\mathop{\arg\max}\limits_{\bm{x}\in\mathbb{DH}^{n},\ \|\bm{x}\|^{2}_{2}=n}\ \bm{x}^{*}\bm{C}\bm{x}$ (4)
@@
By Fact 2.2, the optimum of (4) is attained at the dominant right eigenvector $\bm{u}_{1}$ of $\bm{C}$, normalized so that $\|\bm{u}_{1}\|_{2}^{2}=n$. Recently, Cui and Qi (2024) proposed a power iteration method for computing the dominant eigenvalue and a corresponding eigenvector of a dual quaternion Hermitian matrix, and established linear convergence of the resulting iterates to the dominant eigenpair. The procedure is summarized in Algorithm 1.
@@
1:Input: Hermitian matrix $\bm{C}\in\mathbb{DH}^{n\times n}$ with dominant right eigenvalue $\lambda_{1}$ satisfying $\lambda_{1,\mathrm{st}}\neq 0$, random initial point $\bm{w}^{0}\in\mathbb{DH}^{n}$ such that $(\bm{w}^{0}_{\mathrm{st}})^{*}\bm{u}_{1,\mathrm{st}}\neq 0$.
@@
The mild assumption $\lambda_{1,\mathrm{st}}\neq 0$ and $(\bm{w}^{0}_{\mathrm{st}})^{*}\bm{u}_{1,\mathrm{st}}\neq 0$ ensures that dominant eigenspace exists and the initialization has a nonzero standard part projection onto the dominant eigenspace. If $(\bm{w}^{0}_{\mathrm{st}})^{*}\bm{u}_{1,\mathrm{st}}=0$, then the power iteration cannot amplify (and hence cannot converge to) the dominant eigenspace. Under these assumptions, the standard part iterate remains nonzero, i.e., $\bm{y}^{k}_{\mathrm{st}}\neq\bm{0}$, making the inverse well-defined.
@@
We first define the normalization operator $\mathcal{N}(\cdot)$, which maps a dual quaternion $x\in\mathbb{DH}$ to a unit dual quaternion $\mathcal{N}(x)\in\mathrm{UDQ}$. Let $u:=\mathcal{N}(x)$. If $x_{\mathrm{st}}\neq 0$, then
$u_{\mathrm{st}}=\frac{x_{\mathrm{st}}}{|x_{\mathrm{st}}|},\ u_{\mathcal{I}}=\frac{x_{\mathcal{I}}}{|x_{\mathrm{st}}|}-\frac{x_{\mathrm{st}}}{|x_{\mathrm{st}}|}\mathrm{sc}\left(\frac{x_{\mathrm{st}}^{*}}{|x_{\mathrm{st}}|}\frac{x_{\mathcal{I}}}{|x_{\mathrm{st}}|}\right).$
If $x_{\mathrm{st}}=0$ and $x_{\mathcal{I}}\neq 0$, then
$u_{\mathrm{st}}=\frac{x_{\mathcal{I}}}{|x_{\mathcal{I}}|},\ u_{\mathcal{I}}\text{ is any quaternion satisfying }\mathrm{sc}(x_{\mathcal{I}}^{*}u_{\mathcal{I}})=0.$
@@
$\widetilde{y}_{i}=\begin{cases}\mathcal{N}(y_{i})&\text{if }y_{i}\neq 0,\\ \mathcal{N}(\bm{e}^{*}\bm{y})&\text{otherwise},\end{cases}$
where $\bm{e}\in\mathbb{DH}^{n}$ is any vector satisfying $\bm{e}^{*}\bm{y}\neq 0$. If $\bm{y}=\bm{0}$, we set $\widetilde{\bm{y}}=\bm{1}$. Then $\tilde{\bm{y}}$ is the projection of $\bm{y}$ onto $\mathrm{UDQ}^{n}$, i.e., $\widetilde{\bm{y}}=\Pi(\bm{y})$.
@@
$\widetilde{\bm{x}}:=\Pi(\bm{u}_{1}),$ (5)
which lies in $\mathrm{UDQ}^{n}$ by construction. We next show that $\widetilde{\bm{x}}$ remains close to the ground truth signal $\hat{\bm{x}}$.
@@
For $\widetilde{\bm{x}}$ given by (5), we have
@@
Theorem 2.8 follows from the Lipschitz property of the normalization/projection map in Lemma 2.5. In particular, it implies that $\widetilde{\bm{x}}$ achieves the same order of accuracy as the (generally infeasible) spectral solution $\bm{u}_{1}$ to (4), while ensuring feasibility in $\mathrm{UDQ}^{n}$.
@@ 3 Iterative Estimation Performance
Although Proposition 2.4 and Theorem 2.8 show that both the (infeasible) spectral solution of (4) and its rounded version yield accurate estimates close to the ground truth signal, these estimators alone do not directly solve (3) beyond serving as a starting point. To obtain a fully iterative refinement procedure whose iterates remain feasible, we develop a dual quaternion generalized power method (DQGPM) as the second-stage algorithm.
@@
1:Input: Hermitian matrix $\bm{C}\in\mathbb{DH}^{n\times n}$, initializer $\bm{x}^{0}=\widetilde{\bm{x}}=\Pi(\bm{u}_{1})\in\mathrm{UDQ}^{n}$.
@@
We need to distinguish the algebraic order used in Section 2 from the error metric used in our convergence analysis. While the spectral estimator relies on the lexicographical order of dual numbers to define the dominant eigenpair (strictly prioritizing the standard part), SE(3) synchronization requires bounded errors in both rotation and translation. Therefore, in this section, we analyze the estimation error by controlling the real-valued magnitudes of the standard component $\operatorname{d}_{\mathrm{st}}$ and the dual component $\operatorname{d}_{\mathcal{I}}$ individually. This component-wise analysis is consistent with the algebraic structure of dual numbers, where the convergence of the dual part ($\operatorname{d}_{\mathcal{I}}$) typically depends on the accuracy of the standard part ($\operatorname{d}_{\mathrm{st}}$).
@@
Suppose that the measurement noise satisfies $\|\bm{\Delta}\|_{\mathrm{op},\mathrm{st}}\leq n/350$, $\|\bm{\Delta}\|_{\mathrm{op},\mathcal{I}}\leq n/300$, and the initialization is given by $\bm{x}^{0}=\widetilde{\bm{x}}$. Then the sequence generated by Algorithm 2 satisfies
$\operatorname{d}_{\mathrm{st}}(\bm{x}^{k},\hat{\bm{x}})\leq\left(\frac{1}{10}\right)^{k}\frac{\sqrt{n}}{25}+\frac{700}{53n}\left\|(\bm{\Delta}\hat{\bm{x}})_{\rm st}\right\|_{2}.$ (6)
Moreover, assume that there exists $B\geq 0$ such that
$\max_{k\geq 0}\max_{1\leq i\leq n}|(r_{i}^{k})_{\mathcal{I}}|\leq B,$ (7)
where $\bm{r}^{k}=\hat{\bm{x}}z_{k}$ with $z_{k}\in\operatorname{argmin}_{z\in{\rm UDQ}}\|\bm{x}^{k}-\hat{\bm{x}}z\|_{2}$, the projection inputs are uniformly nondegenerate:
$\min_{k\geq 0}\min_{1\leq i\leq n}\left|\left(\frac{1}{n}Cx^{k}\right)_{i,{\rm st}}\right|\geq\frac{1}{2},$ (8)
and there exists a finite constant $\gamma\geq 0$ such that
$\delta^{k+1}_{\mathcal{I}}\leq\left\|(\bm{x}^{k+1}-\hat{\bm{x}}z_{k})_{\mathcal{I}}\right\|_{2}+\gamma\left\|(\bm{x}^{k+1}-\hat{\bm{x}}z_{k})_{\rm st}\right\|_{2},$ (9)
where $\delta_{\mathcal{I}}^{k+1}:=\|(\bm{x}^{k+1}-\hat{\bm{x}}z_{k+1})_{\mathcal{I}}\|_{2}$. Then for all $k\geq 0$,
$\left|\operatorname{d}_{\mathcal{I}}(\bm{x}^{k},\hat{\bm{x}})\right|\leq\delta_{\mathcal{I}}^{k}\leq\left(\frac{1}{10}\right)^{k}\delta^{0}_{\mathcal{I}}$ (10)
$+\frac{186B+48\gamma+7}{1050}\left[\frac{\sqrt{n}}{25}k\left(\frac{1}{10}\right)^{k-1}+\frac{7000}{477n}\left\|(\bm{\Delta}\hat{\bm{x}})_{\rm st}\right\|_{2}\right]$
$+\frac{10}{9}\left[\frac{7+6B}{1050}\sqrt{n}+\frac{6B+2\gamma}{n}\left\|(\bm{\Delta}\hat{\bm{x}})_{\rm st}\right\|_{2}\right].$